THE HCN-Ca AND HCN-Sr COMPLEXES FORMED ON THE SURFACE OF HELIUM NANODROPLETS: SPECTROSCOPIC PROBES OF SOLVATION DYNAMICS

\underline{G. E. DOUBERLY}, and R E. MILLER, Department of Chemistry, University of North Carolina at Chapel Hill, Chapel Hill, NC 27599.

Complexes of HCN with the heavy alkaline earth metals are formed on the surface of helium nanodroplets. Two bands are observed for the HCN-Ca binary complex, having vibrational band origins that are separated by 1 GHz. The ratio of the intesities of the two bands is found to be strongly droplet size dependent. The species corresponding to the higher frequency band is reminiscent of the HCN-X (X=Na, K, Rb, Cs) complexes, having a large moment of inertia that is droplet size dependent, indicative of a complex bound to the surface of the droplet. The other band is rotationally resolved and is assigned to the solvated complex. IR-IR double resonance spectroscopy of HCN-Ca clearly indicate that vibrational excitation results in a population transfer between the surface bound and solvated species, suggesting that there is a droplet size dependent barrier between the surface and solvated states. A solvated HCN-Sr complex is not observed upon pick-up. However, vibrational excitation results in the solvation of the complex, as observed in the double resonance spectra upon pumping the C-H stretch of the surface bound complex.